OpenStudy (anonymous):

it took peter sixteen more minutes to drive to work today than it did yesterday tomorrow he wants it to take five minutes less than yesterday so that his average drive time for three days is no more than 30 minutes how long did it take peter to get to work each day

OpenStudy (anonymous):

If the time he took yesterday was t, then:

OpenStudy (anonymous):

first day 1 min second day 17 min third day 12 total of 3 days is 30 min

OpenStudy (anonymous):

\[{t+t+16+t-5 \over 3}\le30 \implies 3t+11\le90 \implies t \le26.3\]. So the time he took yesterday has to be less than 26.3 min. If we take \(t=25\), then the time he took today was \(25+16=41\) min, and the time he should take tomorrow is 20 min.

OpenStudy (anonymous):

sorry i missed the average part.... first day 21 min second day 37 min third day 32 this gives daily average of 30 min

OpenStudy (anonymous):

@Anwar your equation formation is wrong... so your answers are wrong

OpenStudy (anonymous):

if tie for yesterday = t time for today = t+16 time for tomorrow = t+16-5 = t+11 so t+ t+16+t+11 ------------ = 30 3 3t + 27 ------- = 30 3 3t + 27 = 90 3t = 90-27 t = 63/3 = 21

OpenStudy (anonymous):

so time for yesterday = 21 min time for today = 21 + 16 = 37 time for tomorrow = t+11 = 21 + 11 = 32

OpenStudy (anonymous):

hey susie check the ans and if satisfied, pls click GOOD ANSWER button for me :})

OpenStudy (anonymous):

The question says five minutes less than yesterday, right?!